OpenStudy (anonymous):

3.Your piggy bank has a total of 46 coins in it; some are dimes and some are quarters. If you have a total of $7.00, how many quarters and how many dimes do you have?

OpenStudy (anonymous):

q + d = 46 .25q+.1d = 7 solve the system of equations

OpenStudy (anonymous):

2q + 2d = 92 5q + 2d = 140 subtract -3q = -48 q = 16 plug q back into original equation q + d = 46 16 + d = 46 d = 30 there are 16 quarters and 30 dimes

OpenStudy (anonymous):

Cool xD

OpenStudy (anonymous):

oh wow I had no clue how to solve that, thank you

OpenStudy (anonymous):

you're welcome. just gotta realize that quarters are .25 dollars each and dimes are .1 dollars each. you you are given the number of coins these problems become a lot easier to realize what your system of equations is. you just have to PRACTICE

OpenStudy (anonymous):

i know, no one likes that word...

OpenStudy (anonymous):

lol that helps a lot actually!

OpenStudy (anonymous):

when you look at my work, realize that i multiply the equations by certain numbers so that the coefficients on my dimes is the same, this allows me to subtract and get rid of the dimes variable, leaving me with just quarters. this allows me to find the number or quarters first, then i can use that number quarters to find the number of dimes

OpenStudy (anonymous):

I understand now, it was just really difficult for me to make them equations and solve them